OpenStudy (anonymous):

Solve the equation. 9x^2+16=0

OpenStudy (lgbasallote):

you're gonna get some complex numbers....

OpenStudy (lgbasallote):

is that how you want it?

OpenStudy (lgbasallote):

i mean if your lesson is complex numebrs?

OpenStudy (anonymous):

\[x^2=16/9\]

OpenStudy (anonymous):

-16/9

OpenStudy (anonymous):

So what @lgbasallote said....

OpenStudy (anonymous):

@beeqay can you show me how you got the answer

OpenStudy (maheshmeghwal9):

Ans is like that:- \[9x^2+16=0\] \[9x^2=0-16=-16\] Therefore,\[x^2=-16\div4\] I think now u can get it:)